a man has 4 pairs of dress pants, 4 dress shirts, and 10 ties. how many different ways can he dress for a job interview
Answer by Edwin McCravy(20054)

He can choose a pair of dress pants 4 ways.

For each of those 4 ways to choose a pair of pants,
he can choose a dress shirt 4 ways.

That's 4*4 = 16 ways he can choose a pair of pants and a shirt.

For each of those 16 ways to choose a pair of pants and a shirt,
he can choose a tie 10 ways.

That's 16*10 = 160 ways he can choose a pair of pants, a shirt,
and a tie. 

Answer and how to get it: 4*4*10 = 160 ways.

Edwin